# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

### 18 Mar 2026 v2

@@ 6 Explicit Lower Bounds for Primes of Cryptographic Interest
The asymptotic $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ should be capable of being made fully explicit by combining explicit Mertens product and $\vartheta(x)$ estimates (e.g., Rosser–Schoenfeld [18], Dusart [19]) with an explicit Linnik-type bound (see Xylouris [20]). We do not pursue the bookkeeping here.
@@
The bound $1/2$ for the infinite product can be sharpened for specific primes: $\prod_{j\geq 1}(1-q^{-j})\geq 1-1/(q-1)$, which exceeds $2/3$ for $q\geq 5$ and is at least $0.99$ for $q\geq 101$. For large NTT primes with $q>2^{30}$, this factor is numerically indistinguishable from 11. Even for the smaller NIST moduli $q=3329$ and $q=8380417$ [21, 22], the elementary bound gives $\prod_{j\geq 1}(1-q^{-j})>0.9996$ and $\prod_{j\geq 1}(1-q^{-j})>0.9999998$, respectively.
We now instantiate the general $\omega(q-1)$-based bound in several parameter regimes relevant to cryptographic applications, including many NTT-friendly moduli.
Proposition 4 (Bounds for primes with bounded $\omega(q-1)$).
Let $q\geq 3$ be a prime. Define
$B(\omega)=P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
Then $c(q)\geq B(\omega(q-1))$. Explicit values:
  1. (a)
    $\omega(q-1)=1$: $B(1)=P_{3}/2\approx 0.280063$.
  2. (b)
    $\omega(q-1)\leq 5$: $B(5)\approx 0.116390$.
  3. (c)
    $\omega(q-1)\leq 10$: $B(10)\approx 0.0884704$.
  4. (d)
    $\omega(q-1)\leq 15$: $B(15)\approx 0.0776918$.
  5. (e)
    $\omega(q-1)\leq 20$: $B(20)\approx 0.0715828$.
Proof.
The bound $c(q)\geq B(\omega(q-1))$ is immediate from Theorem 8. Direct evaluation gives
$B(1)=\frac{P_{3}}{2}\approx 0.280063,\qquad B(5)=\frac{16P_{3}}{77}\approx 0.116390,$
$B(10)\approx 0.0884704,\qquad B(15)\approx 0.0776918,\qquad B(20)\approx 0.0715828.$
This yields the stated bounds. ∎
@@
The claim under $\omega(q-1)\leq 10$ follows from Proposition 4(c), since
@@
By Dusart’s explicit product estimate [19, Theorem 5.9], which we use conservatively with threshold $x\geq 2\,278\,383$,
@@
Let $\Phi$ denote the standard normal distribution function. The shifted-prime analogue of the Erdős–Kac theorem (due to Barban [23]; see also Halberstam [24] and Elliott [25, Ch. 12]) asserts that for every real $z$,
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\left\{q\leq x:q\text{ prime and }\frac{\omega(q-1)-\log\log x}{\sqrt{\log\log x}}\leq z\right\}=\Phi(z).$
Consequently, for every fixed $\delta>0$,
$\frac{1}{\pi(x)}\#\left\{q\leq x:q\text{ prime and }\omega(q-1)>(1+\delta)\log\log x\right\}\longrightarrow 0.$
Thus, for large $x$, almost all primes $q\leq x$ satisfy
$\omega(q-1)\leq(1+\delta)\log\log x,$
which explains why bounds depending only on $\omega(q-1)$ are relevant in cryptographic parameter ranges.
@@
Table 1: The prime moduli appearing in the current NIST standards for ML-KEM [21] and ML-DSA [22]. The fourth column gives the general $\omega(q-1)$-based lower bound for $c(q)$ from Theorem 8, while the fifth gives the sharper lower bound obtained from the exact factorization of q1q-1 together with $P(q)\geq 1-\frac{1}{q-1}$. The exact-factorization bounds imply expected rejection-sampling overheads at most $2.17$ and $3.42$, respectively.
@@
NewHope / FALCON [^1] FALCON is being standardized by NIST as FN-DSA (FIPS 206). [26, 27] $12289$ $^{12}\cdot 3 22$ $3.00$
@@
Table 2: Representative NTT-friendly primes from lattice-based cryptography and FHE implementations, together with $\omega(q-1)$ and the dimension-uniform expected rejection-sampling overhead $1/c(q)$. All overheads are at most $4.00$. The first two rows reproduce the NIST entries from Table 1. The SEAL entry is the default coefficient modulus for poly_modulus_degree = 1024 at the 128-bit security level.
@@
For the 64-bit NTT prime $q=2^{64}-2^{32}+1$ (sometimes called the “Goldilocks prime” in the zero-knowledge community, not to be confused with the Ed448-Goldilocks prime $2^{448}-2^{224}-1$ [35]), used as an NTT-friendly ambient field in implementation-oriented lattice cryptography [34], one has $\omega(q-1)=6$ and $\varphi(q-1)/(q-1)\approx 0.2500$. Since $P(q)>1-10^{-19}$, one obtains $c(q)\approx 0.25$, so roughly a quarter of random $n\times n$ matrices have a primitive-root determinant and the security reduction loses a factor of about $4$.
@@ 7 Implications for PRIM-LWE
However, this is a modulus-dependent nuance rather than a structural obstruction. For any fixed prime $p$, the constant $c(p)>0$ and the reduction remains valid. Moreover, the explicit bounds of Section 6 show that the specific moduli surveyed in Table 2—which cover the NIST standards and several widely deployed FHE primes—all have $c(q)$ concretely bounded away from zero.
@@
We note that the primes typically used in lattice-based cryptography are chosen for NTT compatibility. Concretely, this typically means $q\equiv 1\pmod{2^{t}}$ for moderate $t$. Mathematically, this condition controls only the 2-adic part of q1q-1. It does not by itself prevent q1q-1 from having many odd prime factors, and NTT-friendliness alone does not imply a good lower bound on $c(q)$. What our theorems actually use is the controlled factorization structure of q1q-1, quantified by $\omega(q-1)$ or by the exact factorization. The NTT condition is harmless with respect to the 2-part, since the large power of $2$ contributes only the single Euler factor $(1-\tfrac{1}{2})$ to $\varphi(q-1)/(q-1)$. Among the implementation moduli surveyed in Table 2, every prime satisfies $\omega(q-1)\leq 6$. It is this controlled factorization that yields the good lower bounds on $c(q)$ recorded in Remark 5, and Tables 1 and 2. Thus our results do not by themselves indicate a practical weakness; they quantify how modulus choice affects the reduction loss in the prim-lwe framework.
@@ 8 Concluding Remarks
The proof is short and unconditional, requiring only Dirichlet’s theorem and Mertens’ product formula. We have further established several results that illuminate the finer structure of the density $c(p)$:
@@
    In particular, the decay to zero is extremely slow and is completely quantified.
@@
    The quantity $c(p)$ has a continuous limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; equivalently, the distribution function is strictly increasing on $0,\tfrac{1}{2}]$. This full-support conclusion appears to be new. The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$, as shown by a transport lemma, and is purely singular. Pure singularity was already known in the integer case from Erdős [5]; Tenenbaum [3, p. 263] noted that it carries over to shifted primes.
@@
A natural remaining question is to characterize the limiting law $G$ more precisely—for instance, to understand its moments, its modulus of continuity, or its finer local regularity inside $0,\tfrac{1}{2}]$. This should not be confused with the classical support question for the distribution of $\varphi(n)/n$ over all integers, which remains open.

### 30 Mar 2026 v3

@@ 6 Explicit Lower Bounds for Primes of Cryptographic Interest
The asymptotic $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ can be made fully explicit by combining explicit Mertens product and $\vartheta(x)$ estimates (e.g., Rosser–Schoenfeld [19], Dusart [20]) with an explicit Linnik-type bound (see Xylouris [21]). We do not carry out the explicit computation here.
@@
The bound $1/2$ for the infinite product can be sharpened for specific primes: $\prod_{j\geq 1}(1-q^{-j})\geq 1-1/(q-1)$, which exceeds $2/3$ for $q\geq 5$ and is at least $0.99$ for $q\geq 101$. For large NTT primes with $q>2^{30}$, this factor differs from 11 by less than $10^{-9}$. Even for the smaller NIST moduli $q=3329$ and $q=8380417$ [22, 23], the elementary bound gives $\prod_{j\geq 1}(1-q^{-j})>0.9996$ and $\prod_{j\geq 1}(1-q^{-j})>0.9999998$, respectively.
We now instantiate the general $\omega(q-1)$-based bound in several parameter regimes relevant to cryptographic applications, including many NTT-friendly moduli. Define $B(\omega):=P_{3}\prod_{i=1}^{\omega}(1-1/p_{i})$. By Theorem 11, every prime $q\geq 3$ satisfies $c(q)\geq B(\omega(q-1))$. Direct evaluation gives the following values:
$\omega(q-1)\leq K$ $B(K)$
11 $P_{3}/2\approx 0.280063$
55 $\approx 0.116390$
$10$ $\approx 0.0884704$
$15$ $\approx 0.0776918$
2020 $\approx 0.0715828$
@@
The claim under $\omega(q-1)\leq 10$ follows from the bound $c(q)\geq B(\omega(q-1))$ (Theorem 11), since
@@
By Dusart’s explicit product estimate [20, Theorem 5.9], which we use conservatively with threshold $x\geq 2\,278\,383$,
@@
By the shifted-prime Erdős–Kac theorem (Barban [24]; see also Halberstam [25] and Elliott [26, Ch. 12]), $(\omega(q-1)-\log\log x)/\sqrt{\log\log x}$ converges in distribution to the standard normal as $x\to\infty$ over primes $q\leq x$. In particular, for every fixed $\delta>0$, almost all primes $q\leq x$ satisfy $\omega(q-1)\leq(1+\delta)\log\log x$.
@@
The shifted-prime Erdős–Kac theorem explains why bounds depending only on $\omega(q-1)$ are relevant in cryptographic parameter ranges: for $q\approx 2^{128}$, one has $\log\log q\approx 4.5$, so $\omega(q-1)\leq 10$ holds for the vast majority of such primes.
@@
Table 1: The prime moduli appearing in the current NIST standards for ML-KEM [22] and ML-DSA [23]. The fourth column gives the general $\omega(q-1)$-based lower bound for $c(q)$ from Theorem 11, while the fifth gives the sharper lower bound obtained from the exact factorization of q1q-1 together with $P(q)\geq 1-\frac{1}{q-1}$. The exact-factorization bounds imply dimension-uniform expected rejection-sampling overheads at most $2.17$ and $3.42$, respectively.
@@
NewHope / FALCON [27, 28] $12289$ $^{12}\cdot 3 22$ $3.00$
@@
Table 2: Representative NTT-friendly primes from lattice-based cryptography and FHE implementations, together with $\omega(q-1)$ and the dimension-uniform expected rejection-sampling overhead $1/c(q)$. The displayed overheads are rounded to two decimal places; the exact values differ by less than $10^{-3}$ due to the Euler product $P(q)$. All overheads are at most $4.00$. The first two rows reproduce the NIST entries from Table 1.
@@
For the 64-bit NTT prime $q=2^{64}-2^{32}+1$ (sometimes called the “Goldilocks prime” in the zero-knowledge community, not to be confused with the Ed448-Goldilocks prime $2^{448}-2^{224}-1$ [36]), used as an NTT-friendly ambient field in implementation-oriented lattice cryptography [35], one has $\omega(q-1)=6$ and $\varphi(q-1)/(q-1)\approx 0.2500$. Since $P(q)>1-10^{-19}$, one obtains $c(q)\approx 0.25$, so in every dimension $n$ at least a fraction $\approx 0.25$ of random $n\times n$ matrices have a primitive-root determinant, and the dimension-uniform security-reduction overhead is about $4$.
@@ 7 Implications for PRIM-LWE
However, this is a modulus-dependent artifact rather than a structural obstruction. For any fixed prime $p$, the constant $c(p)>0$ and the reduction remains valid. Moreover, the explicit bounds of Section 6 show that the specific moduli surveyed in Table 2—which cover the NIST standards and several widely deployed FHE primes—all have $c(q)$ concretely bounded away from zero.
@@
We note that the primes typically used in lattice-based cryptography are chosen for NTT compatibility. Concretely, this means $q\equiv 1\pmod{2^{t}}$ for moderate $t$. Mathematically, this condition controls only the 2-adic part of q1q-1. It does not by itself prevent q1q-1 from having many odd prime factors, and NTT-friendliness alone does not imply a good lower bound on $c(q)$. What our theorems actually use is the controlled factorization structure of q1q-1, quantified by $\omega(q-1)$ or by the exact factorization. The NTT condition is harmless with respect to the 2-part, since the large power of $2$ contributes only the single Euler factor $(1-\tfrac{1}{2})$ to $\varphi(q-1)/(q-1)$. Among the implementation moduli surveyed in Table 2, every prime satisfies $\omega(q-1)\leq 6$. It is this controlled factorization that yields the good lower bounds on $c(q)$ recorded in Remark 5, and Tables 1 and 2. Thus our results do not by themselves indicate a practical weakness; they quantify how modulus choice affects the reduction loss in the prim-lwe framework.
@@ 8 Concluding Remarks
The proof is short and unconditional, requiring only Dirichlet’s theorem and Mertens’ product formula. We have further established several results that describe the finer structure of the density $c(p)$:
@@
    In particular, the decay to zero is slow and quantified precisely.
@@
    The quantity $c(p)$ has a continuous limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; equivalently, the distribution function is strictly increasing on $0,\tfrac{1}{2}]$. This full-support conclusion appears to be new. The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$, is purely singular, and, by Proposition 4, the limiting measure has Hausdorff dimension 0. The moments of $G$ are given by convergent Euler products (Theorem 8), and the Mellin transform $\mathbb{E}[X^{s}]$ extends analytically to $\{\operatorname{Re}(s)>0\}$ (Theorem 10). Pure singularity was already known in the integer case from Erdős [5]; Tenenbaum [3, p. 263] noted that it carries over to shifted primes.
@@
Proposition 7 and Corollary 7 determine the right-endpoint behavior of $G$:
$1-G\!\left(\frac{1}{2}-\varepsilon\right)\sim\frac{\kappa}{\log(1/\varepsilon)}\qquad(\varepsilon\downarrow 0),$
so $G$ is not Hölder of any positive order at $1/2$. Proposition 4 shows that the limiting measure has Hausdorff dimension 0. Equivalently, the limiting PRIM-LWE law is a continuous, purely singular, full-support measure on $0,\tfrac{1}{2}]$ that is concentrated on a Borel set of Hausdorff dimension 0. A natural remaining question is therefore to determine a global modulus of continuity of $G$. The squared modulus of the characteristic function admits an explicit Euler product (Proposition 5), but whether
$|\widehat{\mu_{f}}(\tau)|\to 0\qquad\text{or}\qquad\limsup_{|\tau|\to\infty}|\widehat{\mu_{f}}(\tau)|>0$
remains open; by Proposition 6, the latter would give an independent Fourier-analytic proof of pure singularity. This should not be confused with the classical support question for the distribution of $\varphi(n)/n$ over all integers, which remains open.